# Formatting Instructions For NeurIPS 2025

### v1

@@ Appendix C Details on concept steering
For concept discovery, we rely on multi-dimensional concept discovery (MCD) [25]. For every feature vector $\phi$, MCD provides a concept decomposition $\phi=\sum_{i=1}^{n_{c}+1}\phi_{i}$, where $\phi_{i}$ is associated with the concept $i$ (of in total $n_{c}$ concepts), which represents a linear subspace of the feature space, and concept $n_{c}+1$ corresponds to the orthogonal complement of the span of all concept subspaces. The latter is necessary to achieve a complete decomposition not explicitly captured through concept subspaces. For a given feature vector, one can now quantify the contribution $\phi_{i}$ from concept $i$ and visualize its magnitude $|\phi_{i}|_{2}$ across the entire feature map to obtain a spatially resolved concept activation map. One option to align such a coarse concept activation map with the input image is to use bilinear upsampling. This process often leads to rather diffuse concept activation maps. Even though we use MCD for demonstration, this alignment step is a common challenge for most concept-based attribution maps.
@@
Appendix D Spatial composition

### v2

@@ Appendix D Details on concept steering
For concept discovery, we rely on multi-dimensional concept discovery (MCD) (Vielhaben et al., 2023). For every feature vector $\phi$, MCD provides a concept decomposition $\phi=\sum_{i=1}^{n_{c}+1}\phi_{i}$, where $\phi_{i}$ is associated with the concept $i$ (of in total $n_{c}$ concepts), which represents a linear subspace of the feature space, and concept $n_{c}+1$ corresponds to the orthogonal complement of the span of all concept subspaces. The latter is necessary to achieve a complete decomposition not explicitly captured through concept subspaces. For a given feature vector, one can now quantify the contribution $\phi_{i}$ from concept $i$ and visualize its magnitude $|\phi_{i}|_{2}$ across the entire feature map to obtain a spatially resolved concept activation map. One option to align such a coarse concept activation map with the input image is to use bilinear upsampling. This process often leads to rather diffuse concept activation maps. Even though we use MCD for demonstration, this alignment step is a common challenge for most concept-based attribution maps.
@@
D.1 FeatInv-Viz visualization of other classes
To provide a closer look at FeatInv-Viz , we present visualizations of three concepts for the classes Great White Shark, Fly, African Elephant, and Water Ouzel. For the visualization we normalize the respective outputs of Algorithm 1 and threshold it below 0.33 as a binary mask to indicate unaffected regions of the image.
Figure 24: FeatInv-Viz visualization of Great White Shark concepts identified within ConvNeXt’s feature space of the Great White Shark class, which can be associated with shark/water, water and shark. Concept 1 and 2 both seem to show water while concept 1 also highlights parts of the shark.
Figure 25: FeatInv-Viz visualization of Fly concepts identified within ConvNeXt’s feature space of the Fly class, which can be associated with fly’s body, outline/part of fly/leaves and leaves/green. Interestingly, Concept 3 focuses only on the supposedly green background in (see images 3 and 4). The rest of the background does not appear to be part of the concept.
Figure 26: FeatInv-Viz visualization of African Elephant concepts identified within ConvNeXt’s feature space of the African Elephant class, which can be associated with tusks, outline/part of elephant and grass/sky. Unlike with the Great White Shark or Fly class, only the outlines of the elephant concept can be seen here.
Figure 27: FeatInv-Viz visualization of Water Ouzel concepts identified within ConvNeXt’s feature space of the Water Ouzel class, which can be associated with water/sky boundary, outline/part of bird and water/sky. As in the concepts of the African Elephant, only an outline of the bird can be seen in these representations.